# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v4

@@ Abstract
By consolidating scattered knowledge, the literature review provides a comprehensive understanding of the investigated topic. However, reading, conducting, or peer-reviewing review papers generally demands a significant investment of time and effort from researchers. To improve efficiency, this paper aims to provide a thorough review of reviews in the PAMI field from diverse perspectives. First, this paper proposes several article-level, field-normalized, and large language model-empowered bibliometric indicators to evaluate reviews. To facilitate this, a meta-data database dubbed RiPAMI, and a topic dataset are constructed. Second, based on these indicators, the study presents comparative analyses of representative reviews, unveiling the characteristics of publications across various fields, periods, and journals. The newly emerging AI-generated literature reviews are also appraised, and the observed differences suggest that most AI-generated reviews still lag behind human-authored reviews in multiple aspects. Third, we briefly provide a subjective evaluation of representative PAMI reviews and introduce a paper structure-based typology of literature reviews. This typology may improve the clarity and effectiveness for scholars in reading and writing reviews, while also serving as a guide for AI systems in generating well-organized reviews. Finally, this work offers insights into the current challenges of literature reviews and envisions future directions for their development.
@@
literature review, machine intelligence, bibliometric, AI-generated scholar content, automated knowledge synthesis.
@@ 1 Introduction
``The entropy of the universe tends to a maximum."
— Rudolf Clausius
THE entropy of almost all natural and artificial systems in the universe exhibits a continuous increase. This principle also applies to the knowledge system of humanity, which similarly undergoes a perpetual escalation in entropy. Such escalating disorder within the knowledge system can lead to redundant endeavors, adversely affecting the generation of new knowledge. Analogous to the role of gravity in shaping the early universe filled with diverse particles, literature review plays a vital role in the knowledge system by consolidating scattered knowledge. Essentially, a literature review is a scholarly composition that not only demonstrates an understanding of the academic literature on a given topic but also situates this knowledge within a broader context. It compiles the most relevant and significant publications related to a specific research area, thereby offering a comprehensive overview of that field.
Almost all the field boast their own literature reviews, especially for the rapidly developing field of pattern analysis and machine intelligence, including image classification [20, 64, 61, 62], image segmentation [67, 70, 91, 35], object detection [116, 21, 81, 47], natural language processing [69, 36, 63], etc. As reported by the AI Index Report [65], there has been a striking surge in artificial intelligence publications, soaring from 200,000 in 2010 to nearly 500,000 by 2021. This exponential increase has subsequently led to a proliferation of related literature reviews. This trend is clearly illustrated in Fig. 1, which shows a marked increase in the annual publication of reviews, underscoring the growing prevalence of literature reviews in the field.
While comprehensive literature reviews are highly favored, writing and peer-reviewing these survey papers imposes a considerable burden on both authors and reviewers, as it usually involves a broader and more thorough examination of the literature than standard research papers demand. Additionally, readers are eager to ascertain which among the numerous reviews on similar topics is most deserving of their time investment. Moreover, with the advent of AI-generated literature review systems, the landscape is further complicated. The differences between AI-generated and human-authored reviews remain an area for exploration.
@@
Up to now, far too little attention has been paid to the above-mentioned issues. This paper concentrates on reviews within the field of PAMI, with three distinct objectives: (1) to endeavor an automated evaluation of these literature reviews in a relatively scientific manner; (2) to offer a thorough review of existing human-authored and AI-generated literature reviews from diverse perspectives; (3) to pinpoint limitations while providing future insights for literature review.
@@ 1.1 Scope
In this study, we exclusively examine review articles. Grant's comprehensive review [31] concluded that there are at least fourteen review types characterized by methods used, which highlights the fact that different reviews serve different purposes. For instance, a critical review strives to exhibit that the author has carried out extensive research on literature and evaluated it critically. A systematic review aims to systematically explore, assess and integrate research evidence, usually in adherence with review guidelines. However, these types of reviews are relatively rare in the field of PAMI. Given the predominance of literature reviews in this area, we entitle this paper "A Literature Review of Literature Reviews".
Although the quantity of reviews is considerably smaller compared to that of normal papers, it remains impractical to analyze reviews within every field. Therefore, this paper will focus only on reviews in the PAMI field. The reviews being analyzed in this paper are sourced from various publication venues, including conferences, journals, and pre-prints. Reviews from various sources exhibit a diverse range of writing styles, article lengths, and citation counts, providing a representative cross-section of the PAMI field. Considering the recent emergence of deep learning in PAMI, most reviews being investigated were published in the last decade. However, for a comprehensive understanding, we also analyze a few reviews published over ten years ago.
1.2 Contribution
In summary, the main contributions of this survey are detailed as follows:
  • We conduct a comprehensive and detailed review of literature reviews in the field of PAMI. To the best of our knowledge, there have been no scholarly attempts to systematically review the literature reviews.
  • A biography database dubbed RiPAMI is built and released. This database contains meta-data for 2904 literature reviews, including titles, authors, citations and references detail, etc. In addition, we construct a topic key phrase dataset comprising 201 different domain reviews through manual annotation. This dataset can be utilized to evaluate the prompt effectiveness in identifying the paper topic or to conduct further analyses for the community.
  • Subjective evaluations of seminal literature reviews in PAMI are presented. Alongside this, we introduce an organizational structure-based review typology designed to guide both human authors and AI systems in the methodical crafting of literature reviews.
  • We propose the impact and quality indicators for quantitatively evaluating literature reviews. Diverging from traditional, our approaches offer real-time, article-level, and field-normalized quantitative assessment of literature reviews. These automated indicators may help researchers in reading, conducting, and peer-reviewing survey papers, and provide support during the appraisal stage of AI-generated review systems.
  • Based on the proposed evaluation indicators, we thoroughly assess the quality of the selected human-authored literature reviews in the PAMI field. Furthermore, a comparison between human-authored and AI-generated reviews is carried out, which highlights the enduring strengths of human-authored literature reviews and uncovers the limitations of existing AI-generated review systems.
  • We briefly discuss the challenges and provide a preliminary analysis of future directions for literature reviews in the PAMI field, highlighting potential areas for improvement.
All the data and code framework used in this paper are publicly available at https://sway.cloud.microsoft/2TXEuPuNlDKEmC9p.
1.3 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 briefs the evolution of the literature review and details the database constructed for this paper. Section 3 introduces the typology of reviews and provides subjective comments on selected reviews. Section 4 presents the quantitative measures used to gauge the impact and quality of the publications respectively, and compares literature reviews from various journals and fields based on these evaluation indicators. The characteristics of human-authored and AI-generated literature reviews are discussed in Section 5 and a case study is carried out to elucidate what are the advantages and shortcomings of these two types of literature reviews. In Section 6, We further discuss the challenges and future of the literature review. Finally, this paper is concluded in section 7.
2 Background and Setting
2.1 Background of Literature Review
Given the historical evolution and disciplinary variations of literature reviews, it is challenging to discern the very first peer-reviewed literature review in history. However, there is limited research to suggest that contemporary literature reviews can be traced back to the 17th and 18th centuries [37] when the scientific method and the concept of peer review began to take shape. Initially, literature reviews were focused more on assimilating existing knowledge as opposed to following a formalized structure. Scholars used prior works to direct their research and grounded their contributions in an established understanding of the subject. Literature reviews during this period served to prevent duplicated efforts and build upon the work of predecessors.
As scholarly communication advanced and the scientific method became more rigorous, the literature review experienced a notable transformation in both methodology and purpose. Nowadays, a well-executed literature review is more than a basic summary of references; it goes beyond that by organizing and combining the information through analysis and synthesis. Reading a good literature review may benefit both novice and seasoned researchers. For novice researchers, literature review is one of the most important ways to gain knowledge in a specific field. They can quickly learn the basics, grasp fundamental concepts, and gain a comprehensive understanding of major theories. For seasoned researchers, literature reviews may help them keep up with the latest research, identify gaps in the current research field, and avoid duplicating work that has already been done.
2.2 Database
Literature reviews usually cite regular papers in a particular field as references; however, our literature review investigates literature reviews in various fields as the material to be analyzed and synthesized. This section describes how the reviews are selected and provides further details on the construction of the database.
@@ 2.2.1 Data Source
Database Title & Authors Venue Abstract Citations References Source Types Charge
arXiv $\times$ $\times$ $\times$ API-based $\times$
CrossRef Counts Only API-based $\times$
Google Scholar Complete $\times$ Crawler-based $\times$
IEEE Xplore Counts Only $\times$ API-based $\times$
Semantic Scholar Complete API-based $\times$
Web of Science Complete API-based
Scopus Complete API-based
arXiv Data File $\times$ $\times$ $\times$ Snapshots $\times$
CrossRef Data File Counts Only Snapshots $\times$
RiPAMI(Ours) Counts Only Snapshots $\times$
TABLE I: Comparisons between various data source.
@@
Tab. I compares various most commonly used data sources, where ``Counts only'' in the citations column means that the source only records the citation counts, while ``Complete'' signifies that the data source provides a complete list of citations. Unfortunately, none of these sources is perfect. For example, arXiv is a free distribution service and an open-access archive for millions of scholarly articles in various fields. Users can utilize arXiv's APIs to access all of the paper meta-data stored in arXiv databases. Meta-data retrieved from arXiv contains valuable information but fails to query the publication venue, citations, and references. Semantic Scholar seems promising, but it suffers a lower update rate than arXiv and a narrower search scope than Google Scholar. Google Scholar is an online search engine that indexes scholarly literature from a wide range of disciplines and publication formats. It employs automated programs to retrieve files for inclusion in the search results (which are not limited to academic papers, but also include patents, books, etc.). Despite its widespread use, Google Scholar still encounters challenges. Beel [11, 12, 13] argues that Google Scholar places a high weight on citation counts in its ranking algorithm and has therefore been criticized for exacerbating the Matthew effect. Moreover, the citation counts displayed on Google Scholar are subject to manipulation by complete nonsense articles indexed on Google Scholar (e.g. citations from AI-generated pre-print papers published on arXiv should have been ignored). Therefore, a promising engineering solution is to leverage the strengths of the different approaches to overcome the weaknesses of each approach, as demonstrated in this paper.
@@ 2.2.2 Database construction
Figure 2: Schematic representation of the database construction process. From keywords to a SQL-based snapshot RiPAMI, there are three major steps to ensure that the data in RiPAMI is clean, accurate, and reliable. For visual clarity, the polar figure displays only a part of retrieving keywords.
We first detail the database construction approach used for this paper (as illustrated in Fig. 2). To avoid potential copyright and licensing issues, papers are retrieved and downloaded using the arXiv's API. Calls to the API are made by means of HTTP requests to a certain URL. The responses will be cached and stored in an SQL-based database.
As can be seen in Fig. 2, we employed a simple yet effective technique to ensure that the type of paper retrieved is a review paper and highly relevant to the field of pattern recognition. First, we identify 106 keywords based on the scopes of related journals and conferences. These selected keywords include but are not limited to speech recognition, optical character recognition, and self-supervised learning. Next, we conduct a primary retrieving and filter the results following a clear rule, i.e. the title of the paper must contain ``survey'' or ``review'', and the keyword should be included in the abstract. Additionally, we filter out noisy data through both ChatGPT-based and manual double-checking processes.
As mentioned in Sec. 2.2.1, we suggest enriching the meta-data of papers by leveraging a combination of disparate data sources. Considering the potential legal risks of crawling to obtain academic data from Google Scholar, Semantic Scholar API was employed to obtain additional meta-data, such as citation and reference details which are not provided by the arXiv API.
Finally, a total of 2904 eligible papers with meta-data are retrieved and collected. To ensure reproducible experiments and prevent overburdening the server, we construct an SQL-based database dubbed RiPAMI (Reviews in Pattern Analysis and Machine Intelligence, pronounced as /ri:pæmi/). This database stores information related to the paper such as title, abstract, date of publication, venue, citation counts, and reference details, etc.
It should be noted that this paper employs the arXiv API for retrieving literature reviews. As such, there might be potential biases and issues related to incomplete retrieval. That is, papers that are not published on arXiv but meet the criteria will not be included in the database. However, we believe that such a problem is unavoidable. On the one hand, it is difficult for any researcher to guarantee that he or she is able to retrieve the entire relevant literature. On the other hand, most researchers tend to select high-level articles as references to ensure the quality of the literature review, which also leads to the bias issue to some extent. Considering that most articles published on arXiv will also be published subsequently in different conferences and journals, the arXiv API-based retrieval can be regarded as a sample of the full set of literature reviews. We expect that the dataset constructed in this manner shares similar statistical characteristics with the full set of literature reviews.
Given that citation counts vary over time, the date for retrieving citation details is January 1, 2024.
2.2.3 Database Statistics
The database consists of more than 2900 literature reviews from a variety of sources, publication years, and fields. To elucidate the characteristics of the RiPAMI database, we conduct a statistical analysis and plot the result in Fig. 3.
Years of Publication Figure 3 (a) illustrates the distribution of publication years of literature reviews contained in the RiPAMI database. What can be clearly seen in this figure is a growing trend in the number of reviews. This trend has similar characteristics to the one obtained in Fig. 1, e.g. both are steadily increasing and showing a fairly significant increase between 2019 and 2020. This to some extent reflects the consistency between our sample data and the original data in terms of statistics.
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (b), the distribution of literature review references in the RiPAMI database follows a log-normal pattern. The average number of cited references is approximately 140, while the median number is 123.
Number of the Citations We count the citations of papers in RiPAMI and plot them in Fig. 3 (d). A power law distribution of received citations could be found. This phenomenon where a small subset of papers receives the majority of citations is sometimes referred to as the ``Matthew Effect" or the ``Pareto principle", as reported in [68, 17].
3 A Subjective Evaluation of Literature Reviews
3.1 Popular Paper Structure

### v5

@@ Abstract
The rapid advancements in Pattern Analysis and Machine Intelligence (PAMI) have led to an overwhelming expansion of scientific knowledge, spawning numerous literature reviews aimed at collecting and synthesizing fragmented information. This paper presents a thorough analysis of these literature reviews within the PAMI field, and tries to address three core research questions: (1) What are the prevalent structural and statistical characteristics of PAMI literature reviews? (2) What strategies can researchers employ to efficiently navigate the growing corpus of reviews? (3) What are the advantages and limitations of AI-generated reviews compared to human-authored ones? To address the first research question, we begin with a narrative overview to highlight common preferences in composing PAMI reviews, followed by a statistical analysis to quantitatively uncover patterns in these preferences. Our findings reveal several key insights. First, fewer than 20% of PAMI reviews currently comply with PRISMA standards, although this proportion is gradually increasing. Second, there is a moderate positive correlation between the quality of references and the scholarly impact of reviews, emphasizing the importance of reference selection. To further assist researchers in efficiently managing the rapidly growing number of literature reviews, we introduce four novel, real-time, article-level bibliometric indicators that facilitate the screening of numerous reviews. Finally, our comparative analysis reveals that AI-generated reviews currently fall short of human-authored ones in accurately evaluating the academic significance of newly published articles and integrating rich visual elements, which limits their practical utility. Overall, this study provides a deeper understanding of PAMI literature reviews by uncovering key trends, evaluating current practices, and highlighting areas for future improvement.
@@
literature review, PAMI, umbrella review, AI-generated review, bibliometrics
@@ 1 Introduction
THE entropy of almost all natural and artificial systems in the universe exhibits a continuous increase. This principle extends to humanity’s knowledge system, which similarly experiences a steady rise in entropy. Such growing disorder within the knowledge system can lead to redundant efforts, hindering the efficient creation of new knowledge. Analogous to the role of gravity in shaping the early universe filled with diverse particles, the literature review serves as a critical mechanism for organizing dispersed knowledge within the academic sphere. Essentially, a literature review is a scholarly composition that not only demonstrates an understanding of the academic literature on a given topic but also situates this knowledge within a broader context. It compiles the most relevant and significant publications related to a specific research area, thereby offering a comprehensive overview of that field.
Nearly every field features its own body of literature reviews, particularly in the rapidly evolving domain of pattern analysis and machine intelligence. As a foundational technology and research direction for AI, it spans multiple areas such as image classification [21, 77, 73, 74], image segmentation [80, 85, 110, 39], object detection [142, 24, 98, 59], natural language processing [84, 40, 75], etc. As reported by the AI Index Report [78], there has been a striking surge in artificial intelligence publications, soaring from 200,000 in 2010 to nearly 500,000 by 2021. This exponential increase has subsequently led to a proliferation of related literature reviews. This trend is clearly illustrated in Fig. 1, which shows a marked increase in the annual publication of reviews, underscoring the growing prevalence of literature reviews in the field.
@@
The increasing number of literature reviews across various AI subfields has raised several concerns. First, from the authors’ perspective, the characteristics of literature reviews in the PAMI field remain underexplored, including aspects such as structural conventions, statistical patterns, and adherence to established review methodologies. Second, researchers face difficulties navigating the rapidly expanding body of literature reviews, as the sheer volume within the same field makes it harder to identify which reviews are most relevant to their research objectives. Lastly, with the growing prevalence of AI-generated reviews, it becomes crucial to assess their benefits, limitations, and overall reliability compared to traditional human-authored reviews.
1.1 Research Question and Contribution
This paper specifically focuses on reviews within the field of PAMI and tries to answer the following three core Research Questions (RQs):
RQ1: What are the prevalent structural and statistical features shared among literature reviews within the PAMI field? Such information helps researchers develop a more specific understanding of literature reviews, thereby advancing the field’s overall comprehension of the role and significance of these reviews.
RQ2: What strategy can researchers employ to navigate the expanding corpus of literature reviews in PAMI more efficiently? A suitable strategy facilitates readers in navigating a large number of similar reviews and identifying the reviews that best match their interests and objectives.
RQ3: What are the advantages and limitations of AI-generated reviews, and what are the key differences between AI-generated and human-authored literature reviews in PAMI? A comprehensive understanding of AI-generated reviews enables users to benefit from the convenience offered by automated review technologies while minimizing the risk of potential misuse.
To address these questions:
  • We conduct both narrative review and statistical analysis to investigate the writing styles, meta-data distributions, and features of literature reviews in the PAMI field. This information aims to help researchers develop a more in-depth understanding of the nature of literature reviews.
  • We propose four real-time, article-level, and field-normalized bibliometric indicators for the quantitative evaluation of literature reviews. It is believed that the integrated use of these indicators may help researchers make more informed choices when selecting review articles.
  • We evaluate several existing automated review generation systems and conducted a systematic assessment and bibliometric analysis of these AI-generated reviews. The findings highlighted notable differences between AI-generated reviews and those written by humans, particularly in the quality of references, and the richness of visual elements.
In summary, we present a literature review of literature reviews and discuss the common concerns faced by existing reviews in the PAMI field. To the best of our knowledge, there has been limited scholarly attention to addressing these considerations. All the data and code framework used in this paper are publicly available at https://sway.cloud.microsoft/2TXEuPuNlDKEmC9p.
1.2 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 briefly introduces the criteria for literature screening, methods for the RiPAMI construction, and formulas for metric calculations. Section 3 provides a narrative analysis of the structure and content features of the selected reviews. Section 4 presents a statistical analysis of meta-data and content features of reviews in the RiPAMI. Section 5 further discusses how the proposed metrics can assist in efficiently selecting the proper reviews. The characteristics of human-authored versus AI-generated literature reviews are discussed in Sec. 6. In Sec. 7, we explore the challenges and future prospects of literature reviews. Finally, the paper concludes in Sec. 8.
2 Methodology
@@ 2.1 Scope
This study examines all fourteen types of review articles as classified in Grant’s comprehensive typology [33], including narrative review, systematic review, state-of-the-art review, etc. While this research is technically a tertiary or “umbrella” study, we opted for the more commonly understood term “literature review” to improve clarity and acceptance within the PAMI field. In addition, although the quantity of reviews is considerably smaller compared to that of normal papers, it remains impractical to analyze reviews within every field. Therefore, this paper will focus only on reviews in the PAMI field.
2.2 Literature Selection Criteria for Classical Review
To identify the features necessary for statistical analysis, we conduct a classical narrative review to determine the key characteristics of literature reviews in the PAMI field. This review utilizes resources such as Google Scholar, Semantic Scholar, IEEE Xplore, and ScienceDirect, covering publications from major publishers including IEEE, Elsevier, Springer, and others. To ensure that the selected papers are both review articles and highly relevant to the PAMI domain, we adopt a straightforward and effective approach. First, the 106 search keywords are derived from the scopes of prominent journals and conferences in the PAMI field, including terms such as speech recognition, optical character recognition, and self-supervised learning. Second, a filtering criterion is applied: the paper’s title must include either “survey” or “review”, and the specified search keyword must appear in the title or abstract. Finally, to maintain the quality of the included reviews, pre-prints not published in journals or conference proceedings are excluded, ensuring that all selected articles are peer-reviewed. This approach helps to maximize the inclusion of high-quality and relevant literature for subsequent analysis.
2.3 Literature Selection Criteria for Statistical Analysis
The immense number of literature reviews in the PAMI field emphasizes the need for approaches beyond manual selection, analysis, and synthesis. Therefore, to ensure a comprehensive and accurate analysis of literature reviews in the PAMI field, we develop an automated process that simulates the manual selection procedure to construct the RiPAMI database.
@@ 2.3.1 Data Source
TABLE I: Comparisons between various sources: “Counts only” in the citations column means that the source only records the citation counts, while “Complete” signifies that the data source provides a complete list of citations.
Database Title & Authors Venue Abstract Citations References Source Types Free
arXiv $\times$ $\times$ $\times$ API-based
CrossRef Counts Only API-based
Google Scholar Complete $\times$ Crawler-based
IEEE Xplore Counts Only $\times$ API-based
Semantic Scholar Complete API-based
Web of Science Complete API-based $\times$
Scopus Complete API-based $\times$
arXiv Data File $\times$ $\times$ $\times$ Snapshots
CrossRef Data File Counts Only Snapshots
RiPAMI(Ours) Complete Snapshots
@@
Table I compares various most commonly used data sources. However, none of these sources is perfect. For example, arXiv is a free distribution service and an open-access archive for millions of scholarly articles in various fields. However, the arXiv data source fails to provide the publication venue, citations, and references. Semantic Scholar seems promising, but it suffers a lower update rate than arXiv and a narrower search scope than Google Scholar. Google Scholar is an online search engine that indexes scholarly literature from a wide range of disciplines. Although Google Scholar uses powerful automated programs to retrieve files for inclusion in its search results, it still faces challenges in many areas. Beel [9, 10, 11] argues that Google Scholar places a high weight on citation counts in its ranking algorithm and has therefore been criticized for exacerbating the Matthew effect [83]. Moreover, the citation counts displayed on Google Scholar are subject to manipulation by complete nonsense articles indexed on Google Scholar (e.g. citations from AI-generated pre-print papers published on arXiv should have been ignored). Therefore, a promising engineering solution is to leverage the strengths of various approaches to overcome the weaknesses of each approach, as will be detailed in the next subsection.
@@ 2.3.2 Database construction
[…]
Figure 2: A schematic overview of the database construction process. From initial keyword selection to the final SQL-based RiPAMI snapshot, three key steps are implemented to ensure the data in RiPAMI is clean, accurate, and reliable. For simplicity, the polar diagram shows only a subset of the keywords used for retrieval. Since citation counts fluctuate over time, the retrieval date for related information is set to October 2024.
As illustrated in Fig. 2, we implement a structured process to construct the RiPAMI database, ensuring that the selected articles are review papers relevant to the pattern recognition field. The process begins with searching with keywords derived from the scopes of relevant journals and conferences. These keywords are then formatted into query strings for the arXiv API as follows: (ti:“review” OR ti:“survey”) AND (ti:“keyword.lower()” OR abs:“keyword.lower()”). This structure specifically retrieves articles whose titles or abstracts include both the terms “review” or “survey” and the specified keywords. Additionally, we apply regular expressions to check that the abstracts of returned articles include the relevant keywords, ensuring that only appropriate review articles enter the RiPAMI database. To avoid potential copyright and licensing issues, papers are retrieved and downloaded using the arXiv’s API. Calls to the API are made by means of HTTP requests to a certain URL. The responses will be parsed and stored in an SQL-based database. Through this process, we collect a total of 4,106 samples. To further refine our database, we employ a double-checking phase. This involves a GPT-based filter to preliminarily classify papers, followed by a manual validation step to remove any noisy or irrelevant entries. Lastly, we supplement the remaining 3099 entries with additional metadata, including the publication date, citation counts, and reference details. As mentioned in Sec. 2.3.1, we suggest enriching the meta-data of papers by leveraging a combination of disparate data sources. Considering the potential legal risks of crawling to obtain academic data from Google Scholar, Semantic Scholar API was employed to obtain additional meta-data, such as citation and reference details which are not provided by the arXiv API. The final SQL-based snapshot encompasses a wide range of data, including ID, title, publication date, citations, and more, facilitating efficient data retrieval and statistical analysis.
2.4 Bibliometric Indicators for Reviews
Literature reviews reflect the state of their respective fields, highlighting current research priorities and offering a glimpse into emerging trends. Despite their significance, systematic methods for evaluating literature reviews remain underdeveloped. By introducing quantitative tools, bibliometric methods offer a promising solution, enabling researchers to objectively analyze literature reviews and uncover key patterns and trends. Bibliometrics is a research field focused on the use of quantitative methods and statistical analysis to evaluate various aspects of scholarly publications [15]. These methods and analysis, including widely recognized metrics like citation counts and impact factors, offer researchers valuable insights in their daily scientific work, aiding them in making prompt and informed decisions. However, most existing bibliometric methods suffer from several limitations including unfair comparison, misuse, and manipulation as introduced in the “San Francisco Declaration on Research Assessment” (DORA) [30] and the Leiden Manifesto [41]. Notably, many of these metrics were not specifically designed for evaluating literature reviews, further limiting their applicability in this context.
In light of these limitations, we propose two new impact indicators, $TNCSI$ and $IEI$, along with two quality indicators, $RQM$ and $RUI$.
2.4.1 Topic Normalized Citation Success Index (TNCSI)
Most existing bibliometric methods face challenges in conducting cross-disciplinary comparisons due to variations between fields [117, 97, 109, 115]. Although a few methods offer the capability for cross-disciplinary comparison, they either require manual assignment of field-specific keywords or involve complex and costly retrieval to calculate metrics. These drawbacks have, to some extent, hindered individual researchers from adopting these methods, thereby limiting the broader application of the metrics.
To fairly compare the external influence of literature reviews across different fields within an acceptable cost, we propose guiding Large Language Models (LLMs) to generate the topic key phrase and then calculating the success index of the current review within the same topic (see Appendix A.1 for details). The $TNCSI$ estimates the impact of research publications within an LLM-generated topic by normalizing citation counts to a scale between 0 and 1. The formula for calculating the $TNCSI$ is:
[…]
where $\lambda$ is the scale parameter obtained through maximum likelihood estimation. Detailed steps for estimating $\lambda$ and constructing the probability density function are provided in the appendix.
The $TNCSI$ demonstrates favorable mathematical properties and interpretability. First, the $TNCSI$ algorithm employs maximum likelihood estimation to convert the probability mass function into a probability density function. This process ensures that, in theory, the $TNCSI$ differentiates between papers with distinct citation counts, avoiding the assignment of identical values to them. Second, the $TNCSI$ possesses physical significance, representing the probability that a specific paper’s citation count surpasses that of any other paper on the same topic. For example, a paper with a $TNCSI$ of 0.5 means it has more citations than half of the papers within the same topic. Furthermore, the calculation of $TNCSI$ is computationally efficient, as it avoids requiring a complete or precise ranking of all related papers. Instead, it provides a reliable estimation of the probability that the given paper surpasses others in terms of citation count within the same topic.
2.4.2 Impact Evolution Index (IEI)
The $IEI$ is defined as the average slope across $l$ distinct points on a $n=l-1$ degree Bézier curve representing the citation trend over time. The formula for calculating $IEI_{L_{l}}$ is:
$IEI_{L_{l}}=\sum_{a=0}^{l-1}\frac{(y_{i}/x_{i})}{l},$ (2)
where $l$ represents the number of months observed. $x_{i}$ and $y_{i}$ are the components of the tangent vector at the $a$-th point on the Bézier curve, indicating the magnitude along the $x$- and $y$-axes, respectively. More details can be found in Appendix A.2
With the guidance of the $IEI$, researchers may further discern the various literature reviews within the same field that exhibit close citation counts. A higher $IEI$ indicates that an increasing number of studies reference the review, signaling its growing influence and attention. Therefore, in practical applications, researchers may prioritize the study of more promising reviews, as indicated by higher positive $IEI$ values, to mitigate redundancy and improve efficiency.
2.4.3 Reference Quality Measurement (RQM)
The RQM incorporates both the quality and timeliness of references by modifying the Gompertz function, which exhibits a sigmoidal growth pattern—slow at the beginning and end, with a rapid increase in the middle phase:
$RQM=1-e^{-\beta\cdot e^{-(1-ARQ)\cdot S_{mp}}}.$ (3)
In this equation, $\beta$ represents the shift parameter, $ARQ$ denotes the average reference quality, while $S_{mp}$ refers to the median semester count of the reference age, indicating the time from the publication date of the references to the issuance of the review (which will be detailed in the Appendix A.3).
Beyond simply assessing the average impact of references, the $RQM$ integrates references’ timeliness and its influence on review quality. This metric addresses the tendency among authors to rely heavily on classic works, which may no longer fully align with current research trends. Additionally, the adjustment coefficient in $RQM$ is calibrated through statistical and optimization algorithms, tailored to the unique characteristics of each research discipline. This data-driven approach minimizes reliance on heuristic parameter settings and increases RQM’s adaptability across various academic disciplines.
2.4.4 Review Update Index (RUI)
The $RUI$ quantifies the necessity of updating a literature review by combining two indicators: the Coverage Difference Ratio (CDR) and the Review Aging Degree (RAD). The formula for calculating the $RUI$ is:
$RUI=p\cdot CDR+q\cdot RAD.$ (4)
Here, $p$ and $q$ are weighting coefficients, set to 10 and 5, respectively. Detailed definitions and calculations of the $CDR$ and $RAD$ are provided in Appendix A.4.
The proposed $RUI$ integrates both the popularity of the research field and the natural aging of individual reviews. It reflects how reviews require updating over time while taking into account the pace of progress within the field. By balancing these factors, $RUI$ offers real-time insights into how urgently a review needs updating. This urgency may arise from rapid developments in the field or the natural aging of the review’s content. $RUI$ moves beyond the simplistic reliance on publication dates, providing a more contextualized assessment of its timeliness.
2.5 Information Extraction for Reviews
Given that this study aims to identify the common features of reviews within the field, solely relying on manual inspection would be both impractical and resource-intensive. To further investigate the features of reviews in the PAMI field, we employed LLM-based information extraction techniques to analyze the review content and parse the response to structured results. The proposed approach converts unstructured text into structured data, facilitating further analysis and interpretation. Details regarding the information extraction method can be found in Appendix B.
3 Classical Review of Reviews in PAMI
This chapter provides a classic, narrative overview of the PAMI field, aiming to identify common features shared by existing reviews and laying the foundation for the subsequent statistical analysis.
3.1 Popular Structure of Reviews in PAMI